OpenStudy (anonymous):

Can someone help me learn how to write recursive equations. The sequence is 1,5,14,30

OpenStudy (anonymous):

\[f _{_{1}} = 1, f _{2}= 5,\] one possible form is: \[f _{n}=a f _{n-1}+b f _{n-2}\]

OpenStudy (anonymous):

using \[f _{_{1}} = 1, f _{2}= 5, f_{3}=14\] \[f _{n}=a f _{n-1}+b f _{n-2}+c f _{n-3}\] you can write \[f _{n}=2 f _{n-1}+0 f _{n-2}+ 2f _{n-3}\] or \[f _{n}=1 f _{n-1}+3 f _{n-2}+ 1f _{n-3}\]

OpenStudy (anonymous):

I am still confused the options are an = an – 1 + n an = an – 1 + n2 an = an – 1 + 2n an = an – 1 + n3

OpenStudy (anonymous):

I know that the correct answer is an=an-1+n^2 But I am not sure how to get that answer.

OpenStudy (anonymous):

\[a _{n}=a _{n-1} +n \] \[a _{n}=a _{n-1} +n ^{2}\] \[a _{n}=a _{n-1} +2n\] \[a _{n}=a _{n-1} +n ^{3}\] I know the correct answer is \[a _{n}=a _{n-1} +n ^{2}\] I am just not sure how.

OpenStudy (anonymous):

I thought if I wrote it like that it would be easier to see or understand I guess.

OpenStudy (amistre64):

i recall something about find a homogenous and then a particular solution, but different classes on the same subject tend to go about this in different ways

OpenStudy (anonymous):

I am not sure this even matters as far as the science goes. I think that the equation is what matters. I am still not sure how though.

OpenStudy (amistre64):

if you are good at spotting patterns, there is a brute method you can try :) n : 1 2 3 4 ... an : 1 5 14 30 ... there is no common difference or ratio that can be useful, so try some ideas; one is to square the n values and see if you can form a pattern ... which of course works since you already know the answer

OpenStudy (anonymous):